Please answer the following, remember to show all steps: These are all apart of one question


1A) The Ballston Corporation granted Peter 100 incentive stock options earlier this year with a $3 per share exercise price. If Peter elects to exercise the options when the stock is worth $5 a share, how much ordinary income will Peter have to recognize at the time of exercise?      

(A)           $0

(B)           $200

(C)           $300

(D)           $500


1B) How much income does Tracy recognize at grant if she received 100 non-qualified stock options with a $3 exercise price and she exercised the option when the stock was worth $7 a share?

(A)           $0

(B)           $300

(C)           $400

(D)           $700



1C) In order to minimize the tax liability associated with selling shares of stock acquired through incentive stock options, how long should the taxpayer hold the stock?           

                                        

(A)           The earlier of 1 year since date of exercise or 2 years since date of grant.

(B)           The earlier of 2 years since date of exercise or 1 year since date of grant.

(C)           The later of 1 year since date of exercise or 2 years since date of grant.

(D)           The later of 2 years since date of exercise or 1 year since date of grant.







there is no further information to be added. the full question is posted. all of these questions are under the subheading.
Answer #1

1A) (b) $ 200

Ordinary Income = 100 * ($ 5 - $ 3) = $ 200

1B) (c) $ 400

Ordinary Income = 100 * ($ 7 - $ 3) = $ 400

1C) (c) The later of 1 year since date of exercise or 2 years since date of grant.

If the ISO shares are sold after 1 year since date of exercise or 2 years since date of grant date is later, any profit or loss will be a capital gain or loss taxed at the long-term capital gains rates (i.e. Preferential Rates).